Ticket Number: HEA-NOC/20030409-01 Ticket Status: UPDATE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: Internet routing
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20030409 09:20 UTC+1
Ticket Closed: Problem End:
Routing issues within HEAnet.
Some clients may experience difficulties with routing
to outside HEAnet. HEAnet NOC are working on this issue
as a matter of urgency.
20030409 09:20 UTC+1 - DC - Problems noticed with routing.
Routing loop detected - we are working on the problem.
20030409 09:57 UTC+1 - DC - Suboptimal BGP config on Galway
core router suspected as cause of problems. Remedial actions
20030409 10:35 UTC+1 - DC - On advice from our suppliers, we have
focused our investigations on the Dublin core router.
20030409 11:00 UTC+1 - DC - Routing table flushed on Dublin
core routing switch. Checking client connectivity.
Time to Fix:
Unknown at this point.